I need help with (b).

I need help with (b).
Answer #1

Given,

v = (-4.05 x 10^7)t^2 + (2.45 x 10^5)t

a)we know that

a = dv/dt and v = dx/dt => dx = v dt

we will diffrentiate the given equation of v for finding a and integrate for finding x

a = d((-4.05 x 10^7)t^2 + (2.45 x 10^5)t)/dt

a = (-8.1 x 10^7) t + (2.45 x 10^5)

b)for this, a = 0

(-8.1 x 10^7) t + (2.45 x 10^5) = 0

t = 2.45 x 10^5/(8.1 x 10^7) = 3.02 x 10^-3 s

Hence, t = 2.02 x 10^-3 s = 0.00302 s